Layoffs looming. American Products plans to layoff 10% of its employees in its aerospace division and 15% of its employees in its agricultural division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?
Answer by checkley77(12844)

x+y=3,000 or x=3,000-y
.10x+.15y=.12*3000
.10(3,000-y)+.15y=.12*3,000
300-.10y+.15y=360
.05y=360-300
.05y=60
y=60/.05
y=1,200 layed off in the agricultural div.
x=3,000-1,200=1,800 layed off in the aerospace div.
Proof:
.10*1,800+.15*1,200=12*3000
180+180=360
360=360